OpenStudy (anonymous):

express 1.8 as a percentage (i just need the formula for it)

OpenStudy (anonymous):

Multiply it by 100..

OpenStudy (campbell_st):

its 1.8 *100 or you can use the factor that 1 = 100% 0.8 = 80% so 1.8 = 100 + 80 = 180%

OpenStudy (anonymous):

\[x \implies x \times 100 \qquad (percentage)\]

OpenStudy (anonymous):

Where x is any fraction..

OpenStudy (anonymous):

thank you!!

OpenStudy (campbell_st):

hope it helps